OpenStudy (jmansedg):

Find the slope of the line that contains the following pair of points. (1, 2), (5, 2) A. -1 B. 0 C. 4 D. Undefined

OpenStudy (jmansedg):

i think it is c

jimthompson5910 (jim_thompson5910):

Step 1) subtract the y coordinates of each point Step 2) subtract the x coordinates in the same order step 3) divide the differences (y over x). Reduce the fraction as much as possible

OpenStudy (mackenzie6302):

its 4

jimthompson5910 (jim_thompson5910):

Step 1) subtract the y coordinates of each point the y coordinates are 2 and 2, so 2-2 = 0 Step 2) subtract the x coordinates in the same order the x coordinates are 1 and 5, so 5-1 = 4 step 3) divide the differences (y over x). Reduce the fraction as much as possible (y difference)/(x difference) = 0/4 = 0 so the slope of this line is 0. This means we have a flat horizontal line